# DiagnosisArena: Benchmarking Diagnostic Reasoning for Large Language Models

### v3

@@ Abstract
The emergence of groundbreaking large language models capable of performing complex reasoning tasks holds significant promise for addressing various scientific challenges, including those arising in complex clinical scenarios. To enable their safe and effective deployment in real-world healthcare settings, it is urgently necessary to benchmark the diagnostic capabilities of current models systematically. Given the limitations of existing medical benchmarks in evaluating advanced diagnostic reasoning, we present DiagnosisArena, a comprehensive and challenging benchmark designed to rigorously assess professional-level diagnostic competence. DiagnosisArena consists of 1,113 pairs of segmented patient cases and corresponding diagnoses, spanning 28 medical specialties, deriving from clinical case reports published in 10 top-tier medical journals. The benchmark is developed through a meticulous construction pipeline, involving multiple rounds of screening and review by both AI systems and human experts, with thorough checks conducted to prevent data leakage. Our study reveals that even the most advanced reasoning models, o3-mini, o1, and DeepSeek-R1, achieve only 45.82%, 31.09%, and 17.79% accuracy, respectively. This finding highlights a significant generalization bottleneck in current large language models when faced with clinical diagnostic reasoning challenges. Through DiagnosisArena, we aim to drive further advancements in AI’s diagnostic reasoning capabilities, enabling more effective solutions for real-world clinical diagnostic challenges. We provide the benchmark and evaluation tools for further research and development. [^1] https://github.com/SPIRAL-MED/DiagnosisArena
@@ 1 Introduction
Recent advances in the reasoning capabilities of large language models (LLMs) have transformed the landscape for addressing complex scientific problems, such as mathematics and programming OpenAI (2024b); DeepSeek-AI (2025). Using step-by-step problem-solving and iterative refinement processes, LLMs have demonstrated performance that exceeds humans in diverse tasks qin2024o1; Huang et al. (2024); Guan et al. (2025); Chen et al. (2025). Emerging studies have increasingly highlighted the promising potential of integrating these reasoning capabilities into clinical scenarios Nori et al. (2024); Huang et al. (2025); Chen et al. (2024b); Yu et al. (2025); Jiang et al. (2025). However, there remains a considerable gap in assessing the readiness of such models for real-world clinical deployment. Consequently, how to benchmark the clinical diagnostic capabilities of reasoning models has become a focal point for further application.
@@
To quantitatively evaluate the performance of diagnostic outputs, we adopt GPT-4o OpenAI (2024a), one of the most powerful models for knowledge-intensive tasks, as a judge to categorize the relationship between a model’s diagnostic results and the ground truth diagnosis as either “identical”, “relevant”, or “irrelevant”. For each case, we generate five candidate diagnostic outputs and calculate both top-1 and top-5 accuracy scores. We conduct extensive experiments involving proprietary models OpenAI (2024b, a); Anthropic (2024); Team (2024); OpenAI (2025); Deepmind (2025) and open-source models DeepSeek-AI (2025); Bingning Wang et al. (2025); Team (2025b); DeepSeek-AI (2024); Team (2025a), including advanced reasoning models (e.g. o1 and DeepSeek-R1). Results on DiagnosisArena reveal that: (1) Current leading models exhibit notably low performance in diagnosing professional-level clinical cases. Even the most advanced reasoning models, o3-mini, o1, and DeepSeek-R1, achieve only 45.82%, 31.09%, and 17.79%, respectively, while other models struggle to surpass 20%. These findings underscore a substantial gap between existing LLM capabilities and the performance required for professional-level medical diagnostics. (2) We examine potential data leakage effects by examining the publication dates of the original case reports. Our analysis indicates that instances of data leakage within our benchmark are exceedingly rare, as indicated by the absence of significant divergence in model performance for cases published before versus after the training data cut-off date. This observation highlights the necessity of developing more sophisticated paradigms to enhance AI capabilities for addressing challenging clinical problems. (3) By converting the diagnostic task in DiagnosisArena into the multiple-choice format based on model-generated diagnoses as DiagnosisArena-MCQ, we observe a marked increase in model performance, with o1 reaching 61.90%, further suggesting that multiple-choice formats inherently reduce task difficulty and thus fail to accurately reflect the models’ true abilities in addressing complex clinical problems.
@@ 2 Related Work
Reasoning LLMs. Endowing LLMs with reasoning abilities has been regarded as a challenging task. As the overall capabilities of LLMs continue to improve, enhancing their reasoning abilities has increasingly become a focal point of research. Initially, the focus shifted from the few-shot approach to making LLMs mimic the reasoning process to fixed paradigms such as Chain-of-Thought Wei et al. (2022) and ReAct Yao et al. (2023b) to stimulate LLMs’ reasoning abilities. Later, researchers recognized that reasoning is a process that is gradually optimized through trial, error, and correction. Based on this understanding, heuristic search and process-level reward models were introduced. Heuristic search drew inspiration from traditional search algorithms, such as Monte Carlo Tree Search Silver et al. (2017), and expanded the LLM’s single chain of thought into a more complex tree structure of thought, as seen in Yao et al. (2023a); Shinn et al. (2023). On the other hand, the process-level reward model models the reasoning process as a Markov Decision Process, gradually assigning rewards to intermediate steps, guiding the model to generate more accurate chains of thought Lightman et al. (2023); Jiao et al. (2024); Wang et al. (2023a). Then, the advent of OpenAI’s o1 OpenAI (2024b) sparked a wave of research on reasoning. Journey Learning qin2024o1 explores multiple strategies to replicate the o1-like slow-thinking reasoning. O1-coder Zhang et al. (2024) integrates Monte Carlo Tree Search in code-related domains. STILL-2 Min et al. (2024) distills long-form reasoning data, expands potential solution paths, and iteratively optimizes. These efforts culminated in breakthroughs exemplified by DeepSeek’s achievements DeepSeek-AI (2025), which underscore the critical role of reinforcement learning in augmenting reasoning performance. The development of reasoning models is still ongoing…
@@ 3.3 Comparison
Figure 3: Performance of Different Models on DiagnosisArena. (a) The Top-$k$ metric represents the proportion of cases where the correct answer is included among the Top-$k$ predictions generated by the model, ranked in descending order of confidence. The results reveal that while the o3-mini outperforms others, DiagnosisArena remains a significant challenge for all existing models. (b) The MCQ presents the multiple-choice version of DiagnosisArena. A marked increase in model performance can be observed, with o1 reaching 61.90%.
@@ 4.1 Implementation Details
Models. Our experiments include both proprietary and open-source models, as well as domain-specific medical models and general models, such as Baichuan-M1 Bingning Wang et al. (2025), DeepSeek-V3 DeepSeek-AI (2024), GPT-4o OpenAI (2024a), Claude-3.5-Sonnet Anthropic (2024), Qwen2.5-Max Team (2024). Notably, we also focused on assessing models that leverage inference-time scaling, such as open-source QwQ-32B Team (2025b), OpenAI’s o1 OpenAI (2024b), DeepSeek-R1 DeepSeek-AI (2025), Qwen3-235B-A22B Team (2025a), Gemini 2.5 Pro Deepmind (2025), and o3-mini OpenAI (2025). Additionally, we tested both the thinking and non-thinking configurations of Qwen3-235B-A22B, denoted as Qwen3-Thinking (with enable_thinking set to true) and Qwen3-NonThinking (with enable_thinking set to false).
@@ 4.3 Main Results
Figure 3 presents the main evaluation results of LLMs on the DiagnosisArena. Based on our analysis of the experimental data, we draw the following conclusions: (1) Even the most advanced reasoning LLMs struggle with this task. The best-performing model, o3-mini, achieved an accuracy of only 45.82% on DiagnosisArena, while o1 and DeepSeek-R1 performed even worse, with an accuracy as low as 31.09% and 17.79%. This substantial performance gap highlights the significant difficulty of our benchmark and the limitations of current models. (2) Models endowed with explicit reasoning capabilities demonstrate a clear advantage in clinical diagnostic tasks. Notably, even powerful models such as Claude-3.5-Sonnet and Qwen2.5-Max achieve suboptimal performance—below 20% accuracy. In contrast, reasoning-enhanced models, including those with relatively smaller parameter sizes (e.g., QwQ-32B), attain significantly higher accuracy, reaching 25.69%. A particularly illustrative case is DeepSeek-R1, a reasoning-oriented model trained by DeepSeek-V3, which achieves a 13.66% improvement in accuracy over its base model. These findings underscore the critical role of reasoning in clinical diagnosis. (3) Solely prompting LLMs to select the best answer in a multiple-choice format does not accurately reflect their true capabilities in handling clinical tasks. A significant performance improvement is often observed in the multiple-choice setting—for example, o1 achieves 61.90%. Even Baichuan-M1-14B, which scores below 10% in open-ended settings, attains 58.31% in the multiple-choice format. This discrepancy can be attributed to the inherently simplified nature of multiple-choice questions, where predefined options narrow the problem space. In such cases, LLMs can leverage superficial cues or rely on partial knowledge to eliminate incorrect choices and select the most plausible answer. However, this does not constitute a complete deductive reasoning process as required in real-world diagnostic scenarios.

### v4

@@ Abstract
The emergence of groundbreaking large language models capable of performing complex reasoning tasks holds significant promise for addressing various scientific challenges, including those arising in complex clinical scenarios. To enable their safe and effective deployment in real-world healthcare settings, it is urgently necessary to benchmark the diagnostic capabilities of current models systematically. Given the limitations of existing medical benchmarks in evaluating advanced diagnostic reasoning, we present DiagnosisArena, a comprehensive and challenging benchmark designed to rigorously assess professional-level diagnostic competence. DiagnosisArena consists of 1,113 pairs of segmented patient cases and corresponding diagnoses, spanning 28 medical specialties, deriving from clinical case reports published in 10 top-tier medical journals. The benchmark is developed through a meticulous construction pipeline, involving multiple rounds of screening and review by both AI systems and human experts, with thorough checks conducted to prevent data leakage. Our study reveals that even the most advanced reasoning models, o3, o1, and DeepSeek-R1, achieve only 51.12%, 31.09%, and 17.79% accuracy, respectively. This finding highlights a significant generalization bottleneck in current large language models when faced with clinical diagnostic reasoning challenges. Through DiagnosisArena, we aim to drive further advancements in AI’s diagnostic reasoning capabilities, enabling more effective solutions for real-world clinical diagnostic challenges. We provide the benchmark and evaluation tools for further research and development. [^1] https://github.com/SPIRAL-MED/DiagnosisArena
@@ 1 Introduction
Recent advances in the reasoning capabilities of large language models (LLMs) have transformed the landscape for addressing complex scientific problems, such as mathematics and programming OpenAI (2024b); DeepSeek-AI (2025). Using step-by-step problem-solving and iterative refinement processes, LLMs have demonstrated performance that exceeds humans in diverse tasks Qin et al. (2024); Huang et al. (2024); Guan et al. (2025); Chen et al. (2025). Emerging studies have increasingly highlighted the promising potential of integrating these reasoning capabilities into clinical scenarios Nori et al. (2024); Huang et al. (2025); Chen et al. (2024b); Yu et al. (2025); Jiang et al. (2025). However, there remains a considerable gap in assessing the readiness of such models for real-world clinical deployment. Consequently, how to benchmark the clinical diagnostic capabilities of reasoning models has become a focal point for further application.
@@
To quantitatively evaluate the performance of diagnostic outputs, we adopt GPT-4o OpenAI (2024a), one of the most powerful models for knowledge-intensive tasks, as a judge to categorize the relationship between a model’s diagnostic results and the ground truth diagnosis as either “identical”, “relevant”, or “irrelevant”. For each case, we generate five candidate diagnostic outputs and calculate both top-1 and top-5 accuracy scores. We conduct extensive experiments involving proprietary models OpenAI (2024b, a); Anthropic (2024); Team (2024); OpenAI (2025b); Deepmind (2025); OpenAI (2025a) and open-source models DeepSeek-AI (2025); Wang et al. (2025); Team (2025b); DeepSeek-AI (2024); Team (2025a), including advanced reasoning models (e.g. o1 and DeepSeek-R1). Results on DiagnosisArena reveal that: (1) Current leading models exhibit notably low performance in diagnosing professional-level clinical cases. Even the most advanced reasoning models, o3, o1, and DeepSeek-R1, achieve only 51.12%, 31.09%, and 17.79%, respectively, while other models struggle to surpass 20%. These findings underscore a substantial gap between existing LLM capabilities and the performance required for professional-level medical diagnostics. (2) We examine potential data leakage effects by examining the publication dates of the original case reports. Our analysis indicates that instances of data leakage within our benchmark are exceedingly rare, as indicated by the absence of significant divergence in model performance for cases published before versus after the training data cut-off date. This observation highlights the necessity of developing more sophisticated paradigms to enhance AI capabilities for addressing challenging clinical problems. (3) By converting the diagnostic task in DiagnosisArena into the multiple-choice format based on model-generated diagnoses as DiagnosisArena-MCQ, we observe a marked increase in model performance, with o1 reaching 61.90%, further suggesting that multiple-choice formats inherently reduce task difficulty and thus fail to accurately reflect the models’ true abilities in addressing complex clinical problems.
@@ 2 Related Work
Reasoning LLMs. Endowing LLMs with reasoning abilities has been regarded as a challenging task. As the overall capabilities of LLMs continue to improve, enhancing their reasoning abilities has increasingly become a focal point of research. Initially, the focus shifted from the few-shot approach to making LLMs mimic the reasoning process to fixed paradigms such as Chain-of-Thought Wei et al. (2022) and ReAct Yao et al. (2023b) to stimulate LLMs’ reasoning abilities. Later, researchers recognized that reasoning is a process that is gradually optimized through trial, error, and correction. Based on this understanding, heuristic search and process-level reward models were introduced. Heuristic search drew inspiration from traditional search algorithms, such as Monte Carlo Tree Search Silver et al. (2017), and expanded the LLM’s single chain of thought into a more complex tree structure of thought, as seen in Yao et al. (2023a); Shinn et al. (2023). On the other hand, the process-level reward model models the reasoning process as a Markov Decision Process, gradually assigning rewards to intermediate steps, guiding the model to generate more accurate chains of thought Lightman et al. (2023); Jiao et al. (2024); Wang et al. (2023a). Then, the advent of OpenAI’s o1 OpenAI (2024b) sparked a wave of research on reasoning. Journey Learning Qin et al. (2024) explores multiple strategies to replicate the o1-like slow-thinking reasoning. O1-coder Zhang et al. (2024) integrates Monte Carlo Tree Search in code-related domains. STILL-2 Min et al. (2024) distills long-form reasoning data, expands potential solution paths, and iteratively optimizes. These efforts culminated in breakthroughs exemplified by DeepSeek’s achievements DeepSeek-AI (2025), which underscore the critical role of reinforcement learning in augmenting reasoning performance. The development of reasoning models is still ongoing…
@@ 3.3 Comparison
Figure 3: Performance of Different Models on DiagnosisArena. (a) The Top-$k$ metric represents the proportion of cases where the correct answer is included among the Top-$k$ predictions generated by the model, ranked in descending order of confidence. The results reveal that while the o3 outperforms others, DiagnosisArena remains a significant challenge for all existing models. (b) The MCQ presents the multiple-choice version of DiagnosisArena. A marked increase in model performance can be observed, with o1 reaching 61.90%.
@@ 4.1 Implementation Details
Models. Our experiments include both proprietary and open-source models, as well as domain-specific medical models and general models, such as Baichuan-M1 Wang et al. (2025), DeepSeek-V3 DeepSeek-AI (2024), GPT-4o OpenAI (2024a), Claude-3.5-Sonnet Anthropic (2024), Qwen2.5-Max Team (2024). Notably, we also focused on assessing models that leverage inference-time scaling, such as open-source QwQ-32B Team (2025b), OpenAI’s o1 OpenAI (2024b), DeepSeek-R1 DeepSeek-AI (2025), Qwen3-235B-A22B Team (2025a), Gemini 2.5 Pro Deepmind (2025), o3-mini OpenAI (2025b), o4-mini OpenAI (2025a) and o3 OpenAI (2025a). Additionally, we tested both the thinking and non-thinking configurations of Qwen3-235B-A22B, denoted as Qwen3-Thinking (with enable_thinking set to true) and Qwen3-NonThinking (with enable_thinking set to false).
@@ 4.3 Main Results
Figure 3 presents the main evaluation results of LLMs on the DiagnosisArena. Based on our analysis of the experimental data, we draw the following conclusions: (1) Even the most advanced reasoning LLMs struggle with this task. The best-performing model, o3, achieved an accuracy of only 51.12% on DiagnosisArena, while o1 and DeepSeek-R1 performed even worse, with an accuracy as low as 31.09% and 17.79%. This substantial performance gap highlights the significant difficulty of our benchmark and the limitations of current models. (2) Models endowed with explicit reasoning capabilities demonstrate a clear advantage in clinical diagnostic tasks. Notably, even powerful models such as Claude-3.5-Sonnet and Qwen2.5-Max achieve suboptimal performance—below 20% accuracy. In contrast, reasoning-enhanced models, including those with relatively smaller parameter sizes (e.g., QwQ-32B), attain significantly higher accuracy, reaching 25.69%. A particularly illustrative case is DeepSeek-R1, a reasoning-oriented model trained by DeepSeek-V3, which achieves a 13.66% improvement in accuracy over its base model. These findings underscore the critical role of reasoning in clinical diagnosis. (3) Solely prompting LLMs to select the best answer in a multiple-choice format does not accurately reflect their true capabilities in handling clinical tasks. A significant performance improvement is often observed in the multiple-choice setting—for example, o1 achieves 61.90%. Even Baichuan-M1-14B, which scores below 10% in open-ended settings, attains 58.31% in the multiple-choice format. This discrepancy can be attributed to the inherently simplified nature of multiple-choice questions, where predefined options narrow the problem space. In such cases, LLMs can leverage superficial cues or rely on partial knowledge to eliminate incorrect choices and select the most plausible answer. However, this does not constitute a complete deductive reasoning process as required in real-world diagnostic scenarios.